To push a 25.0kg crate up a frictionless incline angled at25.0'to the horizontal , a worker exerts a force of 209N parallelto the incline. As the crate slides 1.5m, how much work is done onthe crate by
a) T he worker's applied force,
b) T he gravitational force on the crate, and
c) The normal force exerted by the incline on the crate?
d) What is the total work done on the crate?
Answer #1

Solution -209x1.5 work done 313.5 J b) gravitational force is acting downward. its component along the incline is, mx gxsin(2
Answer #2
a. work done= 209*1.5=313.5 J
b. gravitational force is acting downward. its component along the incline is m*g*sin(25)
but is acting opposite to the motion of the crate. so work done by this will be negative.
work= -m*g*sin(25)*1.5
=-155.47 J
c.the normal force exerted by the incline on the crate= m*g*cos(25)=222.27
d.total work= 313.5-155.47=158 J